H2SO4+H2O=2H2O+2O2


1
Expert's answer
2021-06-11T01:59:16-0400

Equation=

SO2 + O2 + H2O = H2SO4

Word equation is -: Sulfur dioxide + Oxygen gas + Water = Sulfuric acid

Type of chemical reaction is : For this reaction we have a combination reaction.

Balance Reaction=

2SO2 + O2 + 2H2O = 2H2SO4